where in the a graph do the points Y=-2+-x and y=6-2x meet
Answer by dabanfield(803)

where in the a graph do the points Y=-2+-x and y=6-2x meet
1) y = -2 - x
2) y = 6 - 2x
Based on equation 2 substitute 6 - 2x for y in equation 1. Then we have:
6 - 2x = -2 - x
Solve for x:
x = 8
Substitute 8 for x in equation 2:
y = 6 - 2*8
y = 6 - 16 = -10
The common point is (8,-10).